OpenStudy (anonymous):

simplify i^6 times i^8

OpenStudy (anonymous):

(i^6)(i^8) =i^14 =-1

OpenStudy (anonymous):

but, is i a variable or complex number?

OpenStudy (anonymous):

if its a variable, leave it as i^14

OpenStudy (akshay_budhkar):

\[i^{6+8}=i^{14}=i^{4n+2}=-1\]

OpenStudy (anonymous):

where did the i go? and where did the 4n+2 come from?

OpenStudy (akshay_budhkar):

14= 4 (3) +2 which is of the form 4n +2

OpenStudy (akshay_budhkar):

is "i" complex number?

OpenStudy (anonymous):

wait, ofree, answer the question

OpenStudy (anonymous):

Is i a complex number?

OpenStudy (akshay_budhkar):

or is it a variable?

OpenStudy (akshay_budhkar):

\[x^a \times x^b= x^{a+b}\] Remember that.. i believe it is a variable

OpenStudy (akshay_budhkar):

so your answer is \[i^{14}\]

OpenStudy (anonymous):

What Akshay is doing is showing you how to figure out the system of complex numbers. the root of -1 is called i. Therefore, i^2=-1 and i^4 is 1. 4n+2 refers to this pattern for complex numbers.

OpenStudy (anonymous):

the only options are i, -1, -i, 1 so i^14 is not the answer